Circular No. 6551

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1997W AND 1996bw IN NGC 664 P. Berlind, Fred L. Whipple Observatory (FLWO); and P. Garnavich,

Harvard-Smithsonian Center for Astrophysics (CfA), report their discovery of a second supernova (V = 18.0) in NGC 664 from CCD images obtained with the FLWO 1.2-m telescope on Feb. 1.2 UT by C. Hergenrother. SN 1997W is only 0′′.7 west and 2′′.4 south of SN 1996bw (IAUC 6516), which is currently of comparable brightness (V = 18.2). The new supernova was identified from images obtained for the CfA Supernova Photometry program. No evidence of the new supernova is seen on FLWO images taken on 1996 Dec. 6. Spectra of SN 1997W, obtained with the 1.5-m Tillinghast telescope on Feb. 3.1, show a broad Hα P-Cyg feature on an otherwise smooth blue continuum, indicating that this is a type-II supernova caught at an early stage.

BY DRACONIS R. Mewe and J. Heise, Space Research Organization of the Nether-

lands (SRON); J. M. Muller, SRON and BeppoSAX, Rome; A. Coletta, G. D’Andreta, L. Di Ciolo, G. Spoliti, R. Ricci, and P. Giommi, BeppoSAX, Rome; and L. Piro, Istituto Astrofisica Spaziale, Frascati, report: “For 50 min (during an entire orbit of BeppoSAX) during Jan. 23.863–23.898 UT, the Wide Field Camera 2 (WFC2) onboard BeppoSax detected a transient x-ray source that was not seen over 10 hr during immediately-preceding observations of the same field. The source is located at α = 18^{h}33^{m}.4, δ = +51^{o}46′ (equinox 2000.0; error radius 10′), which is 6^{o} from the center of the WFC2 image. Initially, the source flux was 40 mCrab (8.5 WFC2 counts/s) in the energy band 2–9 keV, decreasing to 14 mCrab towards the end of the observation. At higher energies (up to 28 keV), no flux was found during the entire observation. In the next orbit, the 2–9-keV flux had fallen below the sensitivity limit (over one orbit) of ∼ 10 mCrab. We have identified the source with the dM0e flare star BY Dra (distance 15.6 pc). No other sources with a flux > 10 mCrab are present in x-ray catalogues within 3^{o} of the source. We conclude that a giant stellar flare was detected in x-rays with a remarkably short decay time. The expected countrates for typical M dwarf flare stars like BY Dra are of the order of 0.3–1 WFC2 counts/s (during flare), but the measured countrate (corresponding to an x-ray luminosity of ∼ 2 × 10^{31} erg/s) is exceptionally high — comparable to the strongest flare ever seen with the ROSAT PSPC detector.”

1997 February 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green